You produce 100 units of good X which requires inputs A and B. At the combination of inputs used marginal product of input A is 10, marginal product of input B is 20, price of input A is 4, and price of input B is 4. Do you minimize costs? If not, what do you need to do to minimize the cost?
Answer #1

Answer

No.

Use less of A and more of B inputs

======

the costs minimize when the firm produces at MP per dollar is equal and it uses more of higher MP per dollar and so.

MP per dollar =MP/price of input

MP per dollar of input A=10/4=2.5

MP per dollar of input B=20/4=5

so the firm should increase the input B and reduce the input A up to the MP per dollar from both is equal.